OpenStudy (landon34):

Which of the following equations will have a negative solution? m - 4 2/3 = 7 1/2 m - 7 1/2 = 4 2/3 7 1/2 + m = 4 2/3 -7 1/2 + m = 4 2/3

OpenStudy (landon34):

hi guys could you 2 help?

OpenStudy (just_one_last_goodbye):

#1 m = 37/6

OpenStudy (just_one_last_goodbye):

#2 m = 37/6

OpenStudy (just_one_last_goodbye):

#3 m = -5/6

OpenStudy (landon34):

wow, you are a human calculator.

OpenStudy (just_one_last_goodbye):

#4 37/6

OpenStudy (landon34):

it is the third, thanks.

OpenStudy (just_one_last_goodbye):

So the answer that has a negative solution would be Choice C ^_^

OpenStudy (landon34):

thank you

OpenStudy (just_one_last_goodbye):

thank you and yes I know xD my title says it all

OpenStudy (just_one_last_goodbye):

Please medal me by clicking here ^

OpenStudy (just_one_last_goodbye):

sure

OpenStudy (landon34):

I was drinking my second monster, and my chest started hurting, I told everyone about it, and if I didnt answer, to get help. Everybody started freaking out, and one of the guys got a moderator, the chest pain stopped, and I now have UTI

OpenStudy (landon34):

this happened a little while ago.

OpenStudy (landon34):

i almost got caffiene poisining.

OpenStudy (just_one_last_goodbye):

o-o

OpenStudy (landon34):

I didnt know why I said funny, srry

OpenStudy (landon34):

Im gonna change my profile pic , ill be bak.

OpenStudy (just_one_last_goodbye):

ok ^_^

OpenStudy (landon34):

Im thirsty, im gonna get a energy drink, be bak

OpenStudy (landon34):

ok

OpenStudy (landon34):

heh you thought I was dead

OpenStudy (landon34):

did you?

OpenStudy (landon34):

that joke failed

OpenStudy (landon34):

guys I got another problem, ill tag you both